OpenStudy (anonymous):

can somebody help me with this? 40 + 0.1x = 35 + 0.2x

OpenStudy (anonymous):

you have to simplify and solve for the variable x

OpenStudy (anonymous):

yes what would be the first step?

OpenStudy (anonymous):

yes add -0.2x ro each side

OpenStudy (anonymous):

40-0.1x=35

OpenStudy (anonymous):

40+-0.1x=35 then add 40 to each side 40+-40+-0.1=35+40

OpenStudy (anonymous):

remember you still have the addition in there.

OpenStudy (anonymous):

-0.1x=75

OpenStudy (anonymous):

whats 35+40?

OpenStudy (anonymous):

75

OpenStudy (anonymous):

sorry 35+-40

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

right then -0.1x=-5 then you need to divide each side by -0.1

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

nope x=50

OpenStudy (anonymous):

oh ok my bad haha thanks a lot for the help!:]

OpenStudy (anonymous):

click best response?

OpenStudy (anonymous):

40 + 0.1x = 35 + 0.2x 40-35=0.2x-0.1x 5=0.1x x=5/0.1 x=50